Find x2 in the set of given below

x 12 14 10 12 14 15 12 11 11 10

y 45 60 25 25 70 45 50 50 60 70


1
Expert's answer
2021-11-16T13:51:47-0500
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 & X & Y & XY & X^2 & Y^2\\
 \hline
 & 12 & 45 & 540 & 144 & 2025\\
 & 14 & 60 & 840 & 196 & 3600\\
 & 10 & 25 & 250.8 & 100 & 625\\
 & 12 & 25 & 300 & 144 & 625\\
 & 14 & 70 & 980 & 196 & 4900\\
 & 15 & 45 & 675 & 225 & 2025\\
 & 12 & 50 & 600 & 144 & 2500\\
 & 11 & 50 & 550 & 121 & 2500\\
 & 11 & 60 & 660 & 121 & 3600\\
 & 10 & 70 & 700.6 & 100 & 4900\\
 Sum= & 121 & 500 & 6095 & 1491 & 27300\\
\end{array}""\sum_iX_i=121, \sum_iY_i=500"

"\sum_iX_iY_i=6095,\sum_iX_i=1491, \sum_iY_i=27300"


"\bar{X}=\dfrac{1}{n}\sum_iX_i=\dfrac{121}{10}=12.1"

"\bar{Y}=\dfrac{1}{n}\sum_iY_i=\dfrac{500}{10}=50"

"SS_{XX}=\sum_iX_i^2-\dfrac{1}{n}(\sum_iX_i)^2=1491-\dfrac{(121)^2}{10}"

"=26.9"

"SS_{YY}=\sum_iY_i^2-\dfrac{1}{n}(\sum_iY_i)^2=27300-\dfrac{(500)^2}{10}"

"=2300"

"SS_{XY}=\sum_iX_iY_i-\dfrac{1}{n}(\sum_iX_i)(\sum_iY_i)"

"=6095-\dfrac{121(500)}{10}=45"

"m=slope=\dfrac{SS_{XY}}{SS_{XX}}"

"=\dfrac{45}{26.9}=1.672862"

"n=\bar{Y}-m\bar{X}"

"=50-1.672862(12.1)"

"=29.758364"

Therefore, we find that the regression equation is:


"Y=29.758364+1.672862X"



Correlation coefficient:


"r=\dfrac{SS_{XY}}{\sqrt{SS_{XX}}\sqrt{SS_{YY}}}=\dfrac{45}{\sqrt{26.9}\sqrt{2300}}"

"=0.180914"


No correlation.